2 bodies of different masses are dropped from heights 2m and 8m respectively then the ratio of the time taken by them is _________ m

Grade:10

1 Answers

Arun
6 years ago
Let t1 & t2 be time taken to reach the ground then from the formula, 
h=   1/2 gt^2, 
For first body,   2 =1/2   gt_(1 )^(   2) 
For second body,   8 =1/2   gt_2^(   2) 
 2/8 =(t_1^(  2))/(t_2^(  2) ) 
 t_1/t_2   =  1/2.